OpenStudy (anonymous):

J & B Research is conducting a survey to see if people still live in the state in which they were born. They hired Kimberly, Emily, and Ava to conduct the survey. Kimberly surveyed 400 people with 38% of respondents answering yes. Emily surveyed 364 people with 3 out of 7 respondents answering yes. Ava surveyed 350 people with 0.44 of respondents answering yes. Place the survey takers in order from least number of respondents answering yes to greatest number of respondents answering yes.

OpenStudy (anonymous):

Kim had [400*(.38) = 152] people answer 'yes' Emily had [364*(3/7) = 156] people answer 'yes' Ava had [350*(.44) = 154] people answer 'yes'